Three kilograms of steam undergoes an isothermal process from 27.5 bars and 320°C to 7 bars. Determine (a) the amount of heat transferred, in kJ, and (b) the non-flow work done, in kJ, during the process.


1
Expert's answer
2021-03-16T11:36:14-0400

(a) The amount of heat transferred шs zero for all isothermal processes as the temperature does not change.

(b) The non-flow work done for such a process is


"Q=\Delta E+W=0+W,\\
W=Q.\\\space\\
Q=RT\text{ ln}\frac{p_1}{p_2}=\\\space\\
=8.314·(273.15+320)\text{ ln}\frac{27.5}{7}=\\\space\\
=6750\text{ J, or 6.75 kJ}."